Circular No. 6878

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998an IN UGC 3683 K. Ayani, Bisei Astronomical Observatory; and H. Yamaoka, Kyushu

University, report that a spectrogram of SN 1998an (IAUC 6871, 6877) was obtained with the Bisei 1.01-m telescope on Apr. 15.5 UT. A deep and broad Si II absorption at 613 nm (at the rest frame of UGC 3683) and other features indicate that this is a type-Ia supernova around maximum. The expansion velocity derived from the shift of Si II absorption minimum is ∼ 10 000 km/s.

SUPERNOVA 1998aq IN NGC 3982 Ayani and Yamaoka also report: “A spectrogram of SN 1998aq (IAUC

6875), obtained with the 1.01-m telescope on Apr. 15.6 UT, shows a bluish continuum with broad absorption lines at 612 (prominent; Si II), 575 (Si II), 542 (S II), 485 (Fe II), and 438 nm (Mg II); this indicates that SN 1998aq is of type Ia, ∼ 1 week before maximum. The 575-nm feature may imply that SN 1998aq is of subluminous class, but it resembles that of the typical type-Ia SN 1994D. The expansion velocity derived from the shift of the absorption minimum is 11 000 km/s. The absence of narrow Na D absorption suggests that the interstellar reddening is small. The distance of the host galaxy is comparable to that of the Virgo cluster.”

SAX J1808.4–3658 = XTE J1808–369 W. Heindl, D. Marsden, and P. Blanco, Center for Astrophysics and

Space Sciences, University of California at San Diego, report for the RXTE/ HEXTE team: “Pointed observations of XTE J1808–369 (IAUC 6876, 6877) with the RXTE/HEXTE on Apr. 11.85–11.89 and 13.08–13.10 UT reveal that this unusual source has a very hard, power-law spectrum extending to at least 120 keV. This makes XTE J1808–369 one of the hardest binary x-ray pulsars. No spectral variability was observed between the two observations, and the 15–250-keV spectrum is well-fitted by a power law with photon index of 2.02 ± 0.05. The average flux from both days is 89.5 ± 0.6 mCrab (15–250 keV), which corresponds to 1.3 × 10^{−9} erg cm^{−2} s^{−1} (2–10 keV), in good agreement with the RXTE/PCA value from the Apr. 11.8 observation reported by Marshall et al. (IAUC 6876). Using the orbital model of Chakrabarty and Morgan (IAUC 6877), we confirm pulsations at their period, although with low significance due to limited statistics. The very hard nature of this object further motivates additional observations at hard-x-ray and other wavelengths.”

1998 April 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green